# Existence, degeneracy and stability of ground states by logarithmic Sobolev inequalities

### v1

@@ 1. Introduction
To this end, in Section 2 we recall the notions of energy forms and Dirichlet forms on the Clifford algebra, emphasising the standard von Neumann structure of the latter ([C1], [C2]). We then generalise, in Section 3, the classical tools of convergence in probability and uniform integrability for normal states in noncommutative integration theory (see [Dix], [Se1], [N])) and then we establish that sets of states having uniformly bounded relative entropy, with respect to a reference normal state, are uniformly integrable. In Section 4, we apply this approach to show the stability of the ground state under certain unbounded perturbations of the energy form. In Section 5, we outline how this approach may used to give an alternative, infinitesimal approach based on the logarithmic Sobolev inequality in [Gr2] , to the L. Gross’ result in [Gr1] concerning the existence and uniqueness of the ground state of certain physical Hamiltonians for systems of 1/2 Dirac particles.
@@ 3. Tools of noncommutative integration
3.1. Normal functionals vanishing in probability
@@
3.2. Relative uniform integrability of states
The following generalises the notion of uniform integrability of random variables on a probability space.
@@ 3.2. Relative uniform integrability of states
3.3. Norm convergence to zero, relative vanishing and uniform integrability
The following is a noncommutative version of the classical result by which norm convergence to zero in $L^{1}(X,m)$ is equivalent to convergence to zero in probability plus uniform integrability (see e.g. [DM]).
Theorem 3.6.
A sequence of functionals $\varphi_{n}\in M_{*+}$ is vanishing and relatively uniformly integrable with respect to a faithful $\psi\in M_{*+}$ if and only if it norm converges to zero
Proof.
We already noticed that if the sequence norm converges to zero, then it is vanishing and uniformly integrable with respect to $\psi$. In the opposite direction, consider first the splitting
For $\lambda\in[0,k_{\varepsilon}]$ the second summand can be evaluated as follows
$\begin{split}(\xi_{\varphi_{n}}|\mathbb{E}^{\varphi_{n}}[0,k_{\varepsilon}]\xi_{\varphi_{n}})&=(R_{\varphi_{n}}\xi_{\psi}|\mathbb{E}^{\varphi_{n}}[0,k_{\varepsilon}]R_{\varphi_{n}}\xi_{\psi})\\ &=(\xi_{\psi}|\mathbb{E}^{\varphi_{n}}[0,k_{\varepsilon}]R^{2}_{\varphi_{n}}\mathbb{E}^{\varphi_{n}}[0,k_{\varepsilon}]\xi_{\psi})\\ &\leq k^{2}_{\varepsilon}\cdot(\xi_{\psi}|\mathbb{E}^{\varphi_{n}}[0,k_{\varepsilon}]\mathbb{E}^{\varphi_{n}}[0,k_{\varepsilon}]\xi_{\psi})\\ &\leq k^{2}_{\varepsilon}\cdot(\xi_{\psi}|\mathbb{E}^{\varphi_{n}}[0,k_{\varepsilon}]\xi_{\psi}).\end{split}$
Since $\varphi_{n}\to 0$ relatively to $\psi$, we have
$0\leq\lim_{n}(\xi_{\varphi_{n}}|\mathbb{E}^{\varphi_{n}}[0,k_{\varepsilon}]\xi_{\varphi_{n}})\leq k^{2}_{\varepsilon}\cdot\lim_{n}(\xi_{\psi}|\mathbb{E}^{\varphi_{n}}[0,k_{\varepsilon}]\xi_{\psi})=0$
so that $\lim_{n}\|\varphi_{n}\|_{M_{*+}}\leq\varepsilon$ for all $\varepsilon>0$ and $\lim_{n}\|\varphi_{n}\|_{M_{*+}}=0$. ∎
A variation of the above criterion with a stronger assumption is the following
@@ 3.3. Norm convergence to zero, relative vanishing and uniform integrability
norm converges to zero.
@@
so that $0\leq\lim_{n}(\xi_{\varphi_{n}}|\mathbb{E}^{\varphi_{n}}[0,k_{\varepsilon}]\xi_{\varphi_{n}})\leq k_{\varepsilon}\cdot\lim_{n}(\xi_{\psi}|\xi_{\varphi_{n}})=0$. Thus $\lim_{n}\|\varphi_{n}\|_{M_{*+}}\leq\varepsilon$ for all $\varepsilon>0$ and $\lim_{n}\|\varphi_{n}\|_{M_{*+}}=0$. ∎
@@ 3.4. Uniform integrability of sublevel sets of relative entropy
By a lemma above, $\sup_{\xi\in\Xi_{\mathcal{C}}}\|\xi\|^{2}\leq\|\psi\|_{M_{*}}+\sup_{\varphi\in\mathcal{C}}S(\varphi|\psi)<+\infty$, so that $\Xi_{\mathcal{C}}$ is norm bounded, hence weakly relatively compact in $L^{2}(M)$. If $\xi\in L^{2}_{+}(M)$ is a weak limit point of it in $L^{2}(M)$, there exists a sequence $\xi_{n}\in\Xi_{\mathcal{C}}$ weakly convergent to $\xi$. Since the relative entropy functional $\varphi\mapsto S(\varphi|\psi)$ is uniformly bounded on $\mathcal{C}$, by a lemma above, the sequence $\{\varphi_{\xi_{n}}\in\mathcal{C}:\xi_{n}\in\Xi_{\mathcal{C}}\}$ is a fortiori relatively uniformly integrable. If we had $\lim_{n}(\xi_{\varphi_{n}}|\xi_{\psi})=0$, by a lemma above, we would have $\lim_{n}\|\varphi_{\xi_{n}}\|_{M_{*}}=0$ which is a contradiction since $\lim_{n}\|\varphi_{\xi_{n}}\|_{M_{*}}=\lim_{n}1=1$. Thus, necessarily, $(\xi_{\psi}|\xi)=\lim_{n}(\xi_{\varphi_{n}}|\xi_{\psi})\neq 0$ and since $\xi_{\psi}\in L^{2}_{+}(M,\tau)$ is strictly positive as $\psi$ is faithful, we have $\xi\neq 0$.
@@ 4. Ground states by logarithmic Sobolev inequalities
Hence, we may assume that the sequence is made by positive, unit vectors $\xi_{n}\in\mathcal{F}\cap L^{2}_{+}(M,\tau)$ and consider a weak limit of it $\xi\in L^{2}_{+}(M,\tau)$. Since $\lim_{n}\|\sqrt{H}\xi_{n}-\sqrt{\lambda_{0}}\xi_{n}\|=0$ implies that $\lim_{n}(\sqrt{H}\xi_{n}-\sqrt{\lambda_{0}}\xi_{n})=0$ weakly in $L^{2}(M,\tau)$, we have $\lim_{n}\sqrt{H}\xi_{n}=\lim_{n}(\sqrt{H}\xi_{n}-\sqrt{\lambda_{0}}\cdot\xi_{n})+\lim_{n}\sqrt{\lambda_{0}}\cdot\xi_{n}=\sqrt{\lambda_{0}}\cdot\xi$ weakly in $L^{2}(M,\tau)$. As the operator $(\sqrt{H},\mathcal{F})$ is closed with respect to the norm topology of $L^{2}(M,\tau)$, by Mazur’s Theorem, it is also closed with respect to the weak topology of $L^{2}(M,\tau)$ and we may conclude that $\xi\in D(\sqrt{H})=\mathcal{F}$ and $\sqrt{H}\xi=\lambda_{0}\cdot\xi$ and consequently that $\xi\in D(H)$ and $H\xi=\lambda_{0}\cdot\xi$. What is left to be proved to deduce that $\lambda_{0}$ is an eigenvalue of $H$ is that $\xi\neq 0$. This follows from a lemma above and the fact that by (4.1), the relative entropy $S(\cdot|\psi)$ in uniformly bounded above on the set of states $\mathcal{C}:=\{\varphi_{\xi_{n}}\in M_{*+}:n\in\mathbb{N}\}$

### v2

@@ 1. Introduction
To this end, in Section 2 we recall the notions of energy forms and Dirichlet forms on the Clifford algebra, emphasising the standard von Neumann structure of the latter ([C1], [C2]). We then generalise, in Section 3, the classical tools of convergence in probability and uniform integrability for normal states in noncommutative integration theory (see [Dix], [Se1], [N])) and then we establish that sets of states having uniformly bounded relative entropy, with respect to a reference normal state, are uniformly integrable. In Section 4, we apply this approach to show the stability of the ground state under certain unbounded perturbations of the energy form. In Section 5, we outline how these tools may used to give an alternative, infinitesimal approach based on the logarithmic Sobolev inequality in [Gr2] , to the L. Gross’ result in [Gr1] concerning the existence and uniqueness of the ground state of certain physical Hamiltonians for systems of 1/2 Dirac particles.
@@ 3. Tools of noncommutative integration
3.1. Relative uniform integrability of states
@@
The following generalises uniform integrability of random variables on a probability space.
@@
3.2. Norm convergence to zero and uniform integrability
The following is a noncommutative version of a classical result relating norm convergence in $L^{1}(X,m)$ to uniform integrability (see e.g. [DM]).
@@ 3.2. Norm convergence to zero and uniform integrability
norm converges to zero: $\lim_{n}\|\varphi_{n}\|_{M_{*}}=0$.
@@
so that $0\leq\lim_{n}(\xi_{\varphi_{n}}|\mathbb{E}^{\varphi_{n}}[0,k_{\varepsilon}]\xi_{\varphi_{n}})\leq k_{\varepsilon}\cdot\lim_{n}(\xi_{\psi}|\xi_{\varphi_{n}})=0$. Thus $\lim_{n}\|\varphi_{n}\|_{M_{*+}}\leq\varepsilon$ for all $\varepsilon>0$ and then $\lim_{n}\|\varphi_{n}\|_{M_{*+}}=0$. ∎
@@ 3.3. Uniform integrability of sublevel sets of relative entropy
As in the proof of Theorem 3.9, $\sup_{\xi\in\Xi_{\mathcal{C}}}\|\xi\|^{2}\leq\|\psi\|_{M_{*}}+\sup_{\varphi\in\mathcal{C}}S(\varphi|\psi)<+\infty$, so that $\Xi_{\mathcal{C}}$ is norm bounded, hence weakly relatively compact in $L^{2}(M)$. If $\xi\in L^{2}_{+}(M)$ is a weak limit point of it in $L^{2}(M)$, there exists a sequence $\xi_{n}\in\Xi_{\mathcal{C}}$ weakly convergent to $\xi$. Since the relative entropy functional $\varphi\mapsto S(\varphi|\psi)$ is uniformly bounded on $\mathcal{C}$, by Theorem 3.9, the sequence $\{\varphi_{\xi_{n}}\in\mathcal{C}:\xi_{n}\in\Xi_{\mathcal{C}}\}$ is a fortiori relatively uniformly integrable. If we had $\lim_{n}(\xi_{\varphi_{n}}|\xi_{\psi})=0$, by Theorem 3.6, we would have $\lim_{n}\|\varphi_{\xi_{n}}\|_{M_{*}}=0$ which is a contradiction since $\lim_{n}\|\varphi_{\xi_{n}}\|_{M_{*}}=\lim_{n}1=1$. Thus, necessarily, $(\xi_{\psi}|\xi)=\lim_{n}(\xi_{\varphi_{n}}|\xi_{\psi})\neq 0$ and since $\xi_{\psi}\in L^{2}_{+}(M,\tau)$ is strictly positive as $\psi$ is faithful, we have $\xi\neq 0$.
@@ 4. Ground states by logarithmic Sobolev inequalities
Hence, we may assume that the sequence is made by positive, unit vectors $\xi_{n}\in\mathcal{F}\cap L^{2}_{+}(M,\tau)$ and consider a weak limit of it $\xi\in L^{2}_{+}(M,\tau)$. Since $\lim_{n}\|\sqrt{H}\xi_{n}-\sqrt{\lambda_{0}}\xi_{n}\|=0$ implies that $\lim_{n}(\sqrt{H}\xi_{n}-\sqrt{\lambda_{0}}\xi_{n})=0$ weakly in $L^{2}(M,\tau)$, we have $\lim_{n}\sqrt{H}\xi_{n}=\lim_{n}(\sqrt{H}\xi_{n}-\sqrt{\lambda_{0}}\cdot\xi_{n})+\lim_{n}\sqrt{\lambda_{0}}\cdot\xi_{n}=\sqrt{\lambda_{0}}\cdot\xi$ weakly in $L^{2}(M,\tau)$. As the operator $(\sqrt{H},\mathcal{F})$ is closed with respect to the norm topology of $L^{2}(M,\tau)$, by Mazur’s Theorem, it is also closed with respect to the weak topology of $L^{2}(M,\tau)$ and we may conclude that $\xi\in D(\sqrt{H})=\mathcal{F}$ and $\sqrt{H}\xi=\lambda_{0}\cdot\xi$ and consequently that $\xi\in D(H)$ and $H\xi=\lambda_{0}\cdot\xi$. What is left to be proved to deduce that $\lambda_{0}$ is an eigenvalue of $H$ is that $\xi\neq 0$. This follows from Theorem 3.10 and the fact that by (4.1), the relative entropy $S(\cdot|\psi)$ in uniformly bounded above on the set of states $\mathcal{C}:=\{\varphi_{\xi_{n}}\in M_{*+}:n\in\mathbb{N}\}$